OpenStudy (anonymous):

if y/3+2=9/4 then what is y?

OpenStudy (anonymous):

y=3/4 or .75

OpenStudy (anonymous):

y/3 = 9/4 - 2 y/3 = (9-8)/4 y/3 = 1/4 y=3/4

OpenStudy (anonymous):

y/3+2=9/4 y/3=9/4-2 y=27/4-6 y=27-24/4=3/4